Run Meeting: Feb 22nd, 2019
From Hall A Wiki
Back to:
3:30PM Run Meeting, Counting house meeting room (200C)
RC: Cameron Clarke
Contents
Status
Control access to fix septum power supply
- RC daily update:
- CODA cosmics test with no-beam all Day shift: https://logbooks.jlab.org/entry/3657099
- Monitor turbo: look at scattering chamber vacuum - didn't help, so Jessie removed
- Compton polarimeter chicane magnet:
- Looked at by D.C. power during access
- Problem originates in loose wires for phase advance
- Easily fixed and tuned back and tested by D.C. power
- APEX Septum:
- Jack and company identified the core problem: https://logbooks.jlab.org/entry/3657227
- Problem was not SCRs but rather improper internal power circuit configuration
- The voltage taps were incorrectly configured several weeks ago
- The problem caused imbalance of load inside the transformer that caused the trip when LCW failed recently
- The Septum now works at 1045 Amp set point and has taken several shifts worth of data with no problems
- The Septum readback current is slightly below the set value: https://logbooks.jlab.org/entry/3657382
- LHRS dipole quench protection circuit failure:
- Noticed at 2PM while working on other magnets in the hall
- This circuit is an old part that is not professionally built
- Repair consisted of replacing many parts until finally functionality was restored to LHRS dipole at 8PM: https://logbooks.jlab.org/entry/3657228, https://logbooks.jlab.org/entry/3657232
- Several magnet current regulators fall short ~1% level in readback current, which we fixed by adding ~1% to set value: https://logbooks.jlab.org/entry/3657256, IMPORTANT - eyeballed magnet regulator over-compensation recorded here https://logbooks.jlab.org/entry/3657286
- RHRS Q1 magnet ramps to 900 A but the regulator can't hold it and it fell to 860 A with very poor control - i.e. it is not being regulated: https://logbooks.jlab.org/entry/3657240
- RHRS dipole ramping control failure: https://logbooks.jlab.org/entry/3657200
- The RHRS ramping control failed at 6PM while LHRS dipole work was underway (not tested earlier per Jessie's request: https://logbooks.jlab.org/entry/3657054)
- This is because the RHRS current ramp control is managed by a raspberry pi board which has gone bad
- Under these circumstances the RHRS dipole ramps very quickly, which caused the power supply to fault around 7PM:
- Fixing this will take 2 hours to replace the board
- But we can and successfully have ramped the magnet instead by setting this bad regulator on and manually picking momentum set points in small and safe increments, so no controlled access Friday and continue taking data: https://logbooks.jlab.org/entry/3657237, https://logbooks.jlab.org/entry/3657248, https://logbooks.jlab.org/entry/3657280, (success) https://logbooks.jlab.org/entry/3657316
Optics Sieve and SciFi data
- Beam permit after 8pm magnet repair work concluded: https://logbooks.jlab.org/entry/3657221
- Send 5uA tune beam to dump and verify septum corrector magnets send beam to dump
- Bogdan-Wire horizontal scan to establish proper beam spot positioning: https://logbooks.jlab.org/entry/3657266, complementary to prior run https://logbooks.jlab.org/entry/3650677 (don't trust positions shown in screenshot, use positions that they record and define target with)
- SciFi (in left acceptance) running and timing test: https://logbooks.jlab.org/entry/3657285, followup https://logbooks.jlab.org/entry/3657354
- 1M events of Vertical wire 1 target, 2x2 raster, 0.5uA CW (RHRSTwoArm config)
- Midnight escorted access to put in the sieve slits on both arms (now RHRS dipole has ramped and settled nicely): https://logbooks.jlab.org/entry/3657327
- Hit all 3 vertical wires with sieves in, some concern about V2 wire being missed by small raster was corrected by increasing raster size: https://logbooks.jlab.org/entry/3657345
DAQ optimization
- After sieves were removed in 2AM escorted access Alexandre directed DAQ operations while he tuned the DAQ configurations and timing to optimize production running: https://logbooks.jlab.org/entry/3657379, https://logbooks.jlab.org/entry/3657464
- Several runs were taken with varrying configs and prescale values, see compiled run-list: https://logbooks.jlab.org/entry/3657500
- RHRS Cherenkov FADC timing adjusted to pick up coins: https://logbooks.jlab.org/entry/3657473
- DAQ optimization finished at 5:20 AM: https://logbooks.jlab.org/entry/3657472
Production Data Taking
- Once DAQ optimization completed then 30 uA production beam on Tungsten 2.8% target underway, 1.5kHz coins trigger rate: https://logbooks.jlab.org/entry/3657475, https://logbooks.jlab.org/entry/3657435
- Attempt to go to 40 uA thwarted by increased trip rate, returned to 39, then 36 uA: https://logbooks.jlab.org/entry/3657470, https://logbooks.jlab.org/entry/3657478
- Initial production run 4199 plots: https://logbooks.jlab.org/entry/3657468
- Run 4200 plots: https://logbooks.jlab.org/entry/3657474
- See Owl shift leader summary: https://logbooks.jlab.org/entry/3657518
RHRS FADC Raster Current Problems
- During last night's DAQ timing tuning the RHRS FADC timing was adjusted
- This caused the RHRS raster signal to decrease in quality? https://logbooks.jlab.org/entry/3657695
- Alexandre and Bob Michaels are working on this issue, and John is looking at the data
LHRS S0 HV
- Bogdan requests turning off the HV of the S0 scintillator in the LHRS
Slides/Links
MCC scheduling Whiteboard: http://accboard.acc.jlab.org/
APEX wiki page: https://hallaweb.jlab.org/wiki/index.php/APEX
Bluejeans connection information
To join the Meeting: https://bluejeans.com/256205934
To join via Room System: Video Conferencing System: bjn.vc -or-199.48.152.152 Meeting ID : 256205934
To join via phone : 1) Dial:
+1.408.740.7256 (United States) +1.888.240.2560 (US Toll Free) +1.408.317.9253 (Alternate number) (see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 256205934